OpenStudy (anonymous):

Which of the following systems of equations has no solution? 18x − 6y = 14; 3y = 9x − 7 7x − 4 = 3y; 6y − 3 = 14x 4x + y = 10; 2y = 8x − 2 -5y = x − 6; 3x − 15y = 9

OpenStudy (cwrw238):

the correct choice is the one where one equation is a multiple of the other

OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations 4y = x − 13 2x + 16y = 20 (14, 1/4) (-3, 1) (12, -1/4) (5, -2)